OpenStudy (anonymous):

What is the inverse of f(x)= x over 3 -1? f^-1(y)=?

OpenStudy (anonymous):

\[f(x)=\frac{x}{3}-1\] \[f^{-1}(x)=3(x+1)=3x+3\]\]

OpenStudy (anonymous):

first function says divide by 3 then subtract 1. inverse say add 1 then multiply by 3

OpenStudy (anonymous):

method is write \[y=\frac{x}{3}-1\] and solve for x. i see your question asks for \[f^{-1}(y)\] so we will leave it at that

OpenStudy (anonymous):

\[y=\frac{x}{3}-1\] \[y+1=\frac{x}{3}\] \[3(y+1)=x\] \[f{-1}(y)=3(y+1)=3y+3\]